1. A class tossed coins and recorded 161 heads and 179 tails. What is the experimental probability of heads? Of tails?
2. If you guess the answers at random, what is the probability of getting at least three correct answers on a five question true-or-false quiz.
NOTE: We reviewed both problems in class.

Answer by swincher4391(1107)

P(heads) = 161/(161+179) = 161/340
P(tails) = 179/(161+179) = 179/340
P(at least three correct answers) = P(three)+ P(four) + P(five) = (5 choose 3) * (1/2)^3 *(1/2)^2 + (5 choose 4) *(1/2)^4 * (1/2)^1 + (5 choose 5) * (1/2)^5 * (1/2)^0 = 1/2